40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1418     Entry time: Mon Mar 23 15:50:44 2009
Author: rana 
Type: Configuration 
Category: LSC 
Subject: filters deleted, lsc rebooted 

The LSC time had gone too high. I deleted ~20 filters and rebooted. CPU time came down to 50 usec.

The filters all looked like old trash to me, but its possible they were used.

I didn't delete anything from the DARM, CARM, etc. banks but did from the PD and TM filter banks. You can always go back in time by using the

filter_archive/

ELOG V3.1.3-